OpenStudy (anonymous):

Solve for y: 2/3y + 2 = 8

OpenStudy (bahrom7893):

2y + 6 = 24 2y = 24 - 6 y = 12 - 3 y = 9

OpenStudy (bahrom7893):

and next time use proper notation like: (2/3)y + 2 = 8

OpenStudy (anonymous):

oh ok, sorry about that. and thanks :)